# Unknotting number is not additive under connected sum

### v1

@@ Abstract.
We give the first examples of a pair of knots $K_{1}$,$K_{2}$ in the 3-sphere for which their unknotting numbers satisfy $u(K_{1}\#K_{2})<u(K_{1})+u(K_{2})$ This answers question 1.69(B) from Kirby’s problem list [16], in the negative.
@@ 1. Introduction
The unknotting number (or Gordian number) $u(K)$ of a knot $K$ is one of the most fundamental measures of the complexity of a knot. It is defined as the minimum number of crossing changes, interspersed with isotopy, required to transform a diagram of $K$ to a diagram of the unknot. A wide array of techniques have been employed over the years to compute unknotting numbers, using nearly every new technique that has been introduced into knot theory (see [8], [15], [17], [18], [20], [21], [22], [23], [25], [26] for a selection). Yet to this day there are still nine prime 10-crossing knots whose unknotting numbers remain unknown, a testament to the difficulty of determining this basic invariant. See the Knotinfo site [19] for the most up-to-date list of unresolved knots.
@@
Many knot-theoretic constructions and invariants are known to behave well under the connected sum operation. Polynomial invariants, such as the Alexander [1], Jones [14], and HOMFLY polynomials [12] are multiplicative under connected sum, and many basic numerical invariants are very well-behaved under conneccted sum, such as the 3-genus [29], signature [21], braid index [4], and bridge number [30]. Unknotting number has long been conjectured to be additive under connected sum; this conjecture is implicit in the work of Wendt, in one of the first systematic studies of unknotting number [37]. It is unclear when and where this was first explicitly stated; most references to it call it an ‘old conjecture’. It can be found in the problem list of Gordon [13] from 1977 and in Kirby’s list [16].
@@
Showing that $u(K_{1}\#K_{2})\leq u(K_{1})+u(K_{2})$ is very straighforward; choose diagrams for $K_{1}$ and $K_{2}$ that can be unknotted with $u(K_{1})$ and $u(K_{2})$ crossing changes, respectively, and create the connected sum $K_{1}\#K_{2}$ using these diagrams. Then the same set of $u(K_{1})+u(K_{2})$ crossing changes will transform $K_{1}\#K_{2}$ into (unknot)##\##(unknot) = the unknot. The challenge has always been, essentially, to determine whether one cannot do better than this approach, and so $u(K_{1}\#K_{2})\geq u(K_{1})+u(K_{2})$, as well.
Progress on this conjecture has been slow, however. Perhaps the most significant work in support of it has been Scharlemann’s result [28] that a knot $K$ with $u(K)=1$ is prime; consequently, the connected sum of two non-trivial knots must have unknotting number at least 2. More recently, Yang [38] has shown that the connected sum of $n$ identical knots $K$ whose Alexander polynomial satisfies $\Delta_{K}(t)\neq 1$ has unknotting number at least $n$. Very recently, Applebaum, Blackwell, Davies, Edlich, Juhász, Lackenby, Tomašev, and Zheng [2] employed reinforcment learning to search for counterexamples to this conjecture, and established some interesting consequences for unknotting sequences of connected sums. They also showed how, if the conjecture above were true, one could compute the unknotting numbers of many prime knots.
It was this last perspective - what can we learn if we assume that unknotting number is additive - which inpired the authors of the present work to carry out the computations that led to their discovery. We use the notation $\overline{K}$ to represent the mirror image of the knot $K$. It is an elementary fact that $u(L)=u(\overline{L})$ for every knot $L$.
@@
Unknotting number is not additive under connected sum. In particular, for the knot $K=7_{1}$ (the (2,7)-torus knot), with $u(K)=3$, the connected sum $L=K\#\overline{K}$ satifies $u(L)\leq 5<6=u(K)+u(\overline{K})$.
The additivity of unknotting number appears as Problem 1.69(B) in Kirby’s problem list [16]; this result therefore settles this question (in the negative).
Once we have this one example, it takes little work to find infinitely many more. Because the torus knot $7_{1}=T(2,7)$ appears in a minimal unknotting sequence for each of the torus knots $T(2,2k+1)$, $k\geq 3$, for example, we can show:
@@
See Section 2 for more examples.
@@
Our approach to discovering our main example, described in section 2 below, is modeled on the authors’ previous work settling the Bernhard-Jablan Conjecture [6], and at least partly relies on data generated from that project. We also must credit, and express our immense gratitude to, the authors and developers of Snappy/SnapPea [9]; no part of this project would have been possible without the extensive computations made possible by this software.
@@
The structure of this paper is as follows. In Section 2 we discuss the knots $K=7_{1}$ and $L=K\#\overline{K}$, and give the knot diagrams and crossing changes which establish the failure of additivity for unknotting number. In Section 3 we describe the motivation for our searches and the process by which we carried them out. In Section 4 we discuss the consequences of this example for the computation of unknotting number, and possible future directions. The final section provides python code which will allow the reader to carry out their own verification of the computations used in our main example.
2. Proofs of Theorem 1.2 and Corollary 1.3
@@
The two summands which make up our example in Theorem 1.2 are $K=7_{1}$ and its mirror image $\overline{K}$. $K$ is the $(2,7)$-torus knot. A routine calculation establishes that $K$ has signature $\sigma(K)=6$; a result of Murasugi [21] gives $|\sigma(K)|/2\leq u(K)$, and so $K$ has unknotting number at least 3. Since any 3 crossing changes in its standard 7-crossing diagram results in the unknot, we have $u(K)=3$. Alternatively, one can appeal to the important and more far-reaching result of Kronheimer and Mrowka [17] that the unknotting number of the $(p,q)$-torus knot $T(p,q)$ is $(p-1)(q-1)/2$. It is an elementary result that unknotting number is preserved under mirror image - hold a mirror up to your unknotting sequences - and so $u(\overline{K})=3$, as well. Consequently, $u(K)+u(\overline{K})=3+3=6$.
@@
We will demonstrate, however, that $L$ has unknotting number at most $5$. We do this by giving a diagram of $L$ for which changing two crossings gives (a diagram for) the knot $K_{1}=K14a18636$, in the notation from Snappy. We then show that the knot $K_{2}=K15n81556$ has unknotting number (at most) 2, and has a diagram for which one crossing change yields a diagram for $K_{1}$. Consequently, $L=K\#\overline{K}$ has a diagram for which 2 crossing changes can take us to a diagram of $K_{1}$; an isotopy then takes us to another diagram for $K_{1}$ that admits a crossing change to a diagram of $K_{2}$, which has unknotting number at most $2$. So it takes at most $2+1+2=5$ crossing changes, with isotopies, to transform our diagram of $K\#\overline{K}$ to the unknot. This unknotting sequence therefore has length $5$ (that is, $6$ knots and $5$ crossing changes); consequently, $u(K\#\overline{K})\leq 5$.
@@
The reader can verify in Snappy/Sage [9],[27] that this is a braid representative for $L$ by using the command K=Link(braid_closure=brd) or K=snappy.Link(braid_closure=brd) (calling the above tuple ‘brd’) and employing the simplify() and deconnect_sum() commands to recover the summands $7_{1}$ and $\overline{7_{1}}$ Alternatively, one can start from a diagram for $L=7_{1}\#\overline{7_{1}}$, on the left side of Figure 2, and arrive at the diagram of the knot built by Snappy from the braid word after a few Reidemeister moves, as indicated in Figure 2.
@@
Crossing changes on this diagram can be carried out by changing the signs of entries in the tuple, to create a new braid word representing the knot after crossing change. We have, in the above braid word, underlined the entries relevant to our construction. Recalling that Snappy/python starts labeling entries in a tuple at ‘0’, changing the crossings at indices 0 and 1 yields the braid word
$[\underline{-1},\underline{4},2,3,3,3,2,3,2,2,4,-3,-3,-3,-3,-1,-3,-2,-3,-3]$
which Snappy identifies as belonging to the knot $K_{1}=K14a18636$. A diagram for this knot $K_{1}$ built from the closure of this braid represented by the braid word is given in Figure 3; the two crossing changes made to obtain it from the rightmost diagram in Figure 2 are circled. We also provide a schematic of an isotopy from this 20-crossing projection to a 14-crossing projection of $K_{1}$. Since unknotting number changes by at most one under crossing change, and we can recover $L$ from $K_{1}$ by reversing these $2$ crossing changes, we have $u(L)\leq u(K_{1})+2$
@@
We now demonstrate that $K_{1}$ has unknotting number at most $3$. We do this by making use of a pair of 15-crossing diagrams of the knot $K_{2}=K15n81556$. For this discussion we utilize Dowker-Thistlethwaite (DT) codes [10] for the knots, where, again, each entry of the code representents a crossing in the diagram, and a crossing change can be achieved by changing the sign of an entry of the code. The knot $K_{2}$ has, from the authors’ earlier work on the Bernhard-Jablan conjecture [6], 100 distinct 15-crossing projections. For one of them, with DT code
$[\underline{-4},-16,24,26,18,20,28,22,-2,10,12,30,6,8,14]$,
@@
Figure 4. $K14a18636$ to $K15n81556$
On the other hand, a different diagram for $K_{2}$, with DT code
$[4,12,-24,14,18,2,\underline{20},26,8,10,-28,-30,16,-6,-22]$,
when we change a single crossing, at index 6, yields the code
$[4,12,-24,14,18,2,\underline{-20},26,8,10,-28,-30,16,-6,-22]$,
which Snappy identifies as belonging to the knot $K_{3}=K12n412$. The knot $K_{3}$ has unknotting number 1 (as noted, for example, on the Knotinfo site [19]); and in fact changing the crossing in this DT code, at index 13, yields the code
$[4,12,-24,14,18,2,\underline{-20},26,8,10,-28,-30,16,\underline{6},-22]$,
which is a diagram of the unknot. A sequence of Reidemeister moves from the diagram for $K15n81556$ to the unknot, after making these two crossing changes, is shown in Figure 5. Consequently, we can obtain the unknot starting from $K_{1}=K14a18636$ with three crossing changes; first to $K_{2}=K15n81556$, then to $K_{3}=K12n412$, and finally to the unknot.
Figure 5. $K15n81556$ to the unknot
@@
Extending our main example to an infinite class of examples takes little extra work. A knot $K_{1}$ is called Gordian adjacent to the knot $K_{2}$ [11] if $u(K_{2})>u(K_{1})$ and there are $u(K_{2})-u(K_{1})$ crossing changes on a diagram for $K_{2}$ that results in a diagram for $K_{1}$. Put differently, $K_{2}$ possesses a minimal unknotting sequence which contains the knot $K_{1}$. For any pair of knots $L_{1},L_{2}$ that the knot $7_{1}$ is Gordian adjacent to, we can build an unknotting sequence for $L_{1}\#\overline{L_{2}}$ using fewer than $u(L_{1})+u(L_{2})$ crossing changes. Starting with $L_{1}\#\overline{L_{2}}$, applying $u(L_{1})-3=u(L_{1})-u(7_{1})$ crossing changes to the $L_{1}$-summand, and $u(L_{2})-3$ crossing changes to the $\overline{L_{2}}$ summand will yield a diagram for $7_{1}\#\overline{7_{1}}$, and this knot can be unknotted with $5$ crossing changes (after isotopy). Conequently, $(u(L_{1})-3)+(u(L_{2})-3)+5=u(L_{1})+u(L_{2})-1$ crossing changes suffice to unknot $L_{1}\#\overline{L_{2}}$, and so $u(L_{1}\#\overline{L_{2}})\leq u(L_{1})+u(L_{2})-1<u(L_{1})+u(L_{2})$.
One family of knots satisfying this condition is the set of torus knots $T(2,2k+1)$ with $k\geq 3$. These knots have unknotting number $(2-1)((2k+1)-1)/2=k$, and changing any $k-3$ crossings in their standard diagrams will yield $T(2,7)=7_{1}$ (since changing a crossing in the standard diagram for $T(2,N)$ will, after a Reidemeister II move, yield the standard diagram for $T(2,N-2)$, and we are then done by induction). This yields Corollary 1.3. $\square$
Even more, Feller [11] shows that for the torus knots $T(a,b)$ with $gcd(a,b)=1$ and $a\geq 2$, $b\geq 7$ the knot $T(2,7)=7_{1}$ is Gordian adjacent to each of these knots, and, in addition, $T(2,7)$ is Gordian adjacent to the knot $T(3,5)$. All of these knots can therefore also be used to build a connected sum that fails additivity.
@@
For all integers $a_{1},a_{2}\geq 2$ and $b_{1},b_{2}\geq 7$ (or $b_{i}\geq 5$ if $a_{i}=3$) with $\gcd(a_{i},b_{i})=1$ for $i=1,2$, we have
$u(T(a_{1},b_{1})\#\overline{T(a_{2},b_{2})})<u(T(a_{1},b_{1}))+u(T(a_{2},b_{2}))\ .$
For an example which is not a torus knot, we can use the knot $10_{139}$ (which, in fact, was a summand of the first example we found in proving Theorem 1.2; see Section 3 below). The knot $L_{1}=10_{139}$ is a non-alternating hyperbolic knot; it was shown by Kawamura [15] that $u(L_{1})=4$. In fact, Kawamura showed, even more, that $L_{1}$ has smooth 4-genus $g_{4}(L_{1})=4$ (and $g_{4}(L)\leq u(L)$ for every knot $L$ [21]); the 4-genus of $L_{1}$ was also established by different means (using their tau invariant) by Ozsváth and Szabó [24]. One DT code for $10_{139}$ is given by
@@
which belongs to the knot $7_{1}$. As a result, an unknotting sequence for $7_{1}$ (with $3$ crossing changes) can be augmented with this crossing change to yield an unknotting sequence for $10_{139}$ with $4$ crossing changes, that is, a minimal unknotting sequence. Consequently, $10_{139}$ can play the role of $L_{1}$ (and $L_{2}$) in the above discussion, yielding:

### v2

@@ Abstract.
We give the first examples of a pair of knots $K_{1}$$K_{2}$ in the 3-sphere for which their unknotting numbers satisfy $u(K_{1}\#K_{2})<u(K_{1})+u(K_{2})$ . This answers question 1.69(B) from Kirby’s problem list [24], in the negative.
@@ 1. Introduction
For knots $K$ in the 3-sphere $S^{3}$, the unknotting number (or Gordian number) $u(K)$ of $K$ is one of the most fundamental measures of the complexity of the knot. It is defined as the minimum number of crossing changes, interspersed with isotopy, required to transform a diagram of $K$ to a diagram of the unknot. A wide array of techniques have been employed over the years to compute unknotting numbers, using nearly every new technique that has been introduced into knot theory (see [13], [22], [25], [26], [28], [29], [31], [32], [35], [36], [44] for a selection). Yet to this day there are still nine prime 10-crossing knots whose unknotting numbers remain unknown, a testament to the difficulty of determining this basic invariant. See the Knotinfo site [27] for the most up-to-date list of unresolved unknotting numbers.
@@
Many knot-theoretic constructions and invariants are known to behave well under the connected sum operation. Polynomial invariants, such as the Alexander [3], Jones [20], and HOMFLY polynomials [17] are multiplicative under connected sum, and many basic numerical invariants are very well-behaved under connected sum, such as the 3-genus [39], signature [29], braid index [9], and bridge number [40]. Unknotting number has long been conjectured to be additive under connected sum; this conjecture is implicit in the work of Wendt, in one of the first systematic studies of unknotting number [50]. It is unclear when and where this was first explicitly stated; most references to it call it an ‘old conjecture’. It can be found in the problem list of Gordon [19] from 1977 and in Kirby’s list [24].
@@
The inequality $u(K_{1}\#K_{2})\leq u(K_{1})+u(K_{2})$ is always true, and is very straightforward; choose diagrams for $K_{1}$ and $K_{2}$ that can be unknotted with $u(K_{1})$ and $u(K_{2})$ crossing changes, respectively, and create the connected sum $K_{1}\#K_{2}$ using these diagrams. Then the same set of $u(K_{1})+u(K_{2})$ crossing changes will transform $K_{1}\#K_{2}$ into (unknot)#\#(unknot) = the unknot. The challenge has always been, essentially, to determine that one cannot do better than this approach, and so $u(K_{1}\#K_{2})\geq u(K_{1})+u(K_{2})$, as well.
Progress on this conjecture has been slow, however. Several authors have found families $\mathcal{K}$} of knots for which additivity holds for pairs of knots chosen within the family [25],[5]. Perhaps the most significant work in support of the conjecture has been Scharlemann’s result [38] that a knot $K$ with $u(K)=1$ is prime; consequently, the connected sum of two non-trivial knots must have unknotting number at least 2. So unknotting number is additive, when both summands have unknotting number 1. It is also well-known that if two knots $K_{1}$ and $K_{2}$ satisfy $|\sigma(K_{i})|=2u(K_{i})$ for each $i$, where $\sigma(K)$ is the signature of $K$, and $\sigma(K_{1})$ and $\sigma(K_{2})$ have the same sign, then $u(K_{1}\#K_{2})=u(K_{1})+u(K_{2})$. This is a consequence of the inequality $|\sigma(K)|\leq 2u(K)$ [29] and the additivity of signature under connected sum. Since $|\sigma(K_{i})|=2u(K_{i})$ is actually quite common among low crossing number knots (see, e.g., [27]), this offers strong support to the conjecture.
More recently, Yang [52] has shown that the connected sum of $n$ identical knots $K$ whose Alexander polynomial satisfies $\Delta_{K}(t)\neq 1$ has unknotting number at least $n$. Very recently, Applebaum, Blackwell, Davies, Edlich, Juhász, Lackenby, Tomašev, and Zheng [4] employed reinforcment learning to search for counterexamples to Conjecture 1.1, and established some interesting consequences for unknotting sequences of connected sums. They also showed how, if Conjecture 1.1 were true, one could use this to compute the unknotting numbers of many prime knots.
It was this last perspective - what can we learn if we assume that unknotting number is additive - which inspired the authors of the present work to carry out the computations that led to their discovery. In the following we use the notation $\overline{K}$ to represent the mirror image of the knot $K$. It is an elementary fact that $u(K)=u(\overline{K})$ for every knot $K$.
@@
Unknotting number is not additive under connected sum. In particular, for the knot $K=7_{1}$ (the (2,7)-torus knot), with $u(K)=3$, the connected sum $L=K\#\overline{K}$ satisfies $u(L)\leq 5<6=u(K)+u(\overline{K})$.
The additivity of unknotting number appears as Problem 1.69(B) in Kirby’s problem list [24]; Theorem 1.2 therefore settles this question (in the negative).
Once we have this one example, it takes little work to find infinitely many more. The torus knot $7_{1}=T(2,7)$ appears in a minimal unknotting sequence for each of the torus knots $T(2,2k+1)$, $k\geq 3$; see Section 2. From this we can show:
@@
Section 2 describes further examples.
The knot group $\pi(K)=\pi_{1}(S^{2}\setminus K)$, the fundamental group of the complement of the knot in the $3$-sphere, is a powerful tool for the study of a knot. For many knots, it determines the knot up to isotopy and mirroring [51],[18]. However, the knot group does not detect mirroring, and, in particular, $\pi(K\#L)\cong\pi(K\#\overline{L})$ for any knots $K$ and $L$ (see, for example, [23, Corollary 6.1.10]). Since $\sigma(7_{1}\#7_{1})=12$, and so $u(7_{1}\#7_{1})=6$, while $u(7_{1}\#\overline{7_{1}})\leq 5$, we have
Corollary 1.4.
The knot group $\pi(K)$ of a knot $K$ does not determine the unknotting number of $K$.
Using work of Baader [6] we can show that, using the examples of Corollary 1.3, there are infinitely many non-torus knots for which the theorem holds, as well.
Corollary 1.5.
For every $n\geq 4$ there are infinitely many knots $K_{i}$ with $u(K_{i})=n$ and $u(K_{i}\#\overline{K_{j}})<u(K_{i})+u(\overline{K_{j}})$ for every $i$ and $j$.
By taking connected sums of knots idenified in the above corollaries, we can make the additivity gap arbitrarily large.
Corollary 1.6.
For every $n\geq 1$ there are infinitely many knots $K_{i}$ with $u(K_{i}\#\overline{K_{i}})\leq u(K_{i})+u(\overline{K_{i}})-n$.
Further methods for building knots with large gap are discussed in Section 4.
@@
Our approach to discovering our main example, described in section 3 below, is modeled on the authors’ previous work settling the Bernhard-Jablan Unknotting Conjecture [11], and at least partly relies on data generated from that project. We also must credit, and express our immense gratitude to, the authors and developers of SnapPy/SnapPea [14]; no part of this project would have been possible without the extensive computations made possible by this software.
@@
The structure of this paper is as follows. In Section 2 we discuss the knots $K=7_{1}$ and $L=K\#\overline{K}$, and give the knot diagrams and crossing changes which establish the failure of additivity for unknotting number. We also show how to use this result to find many more examples where additivity fails. In Section 3 we describe the motivation for our searches, including work on a related conjecture, on the invariance of unknotting number under knot mutation (which is Problem 1.69(C) of [24]), and outline the process by which we carried the searches out. In Section 4 we discuss the consequences of this example for the computation of unknotting number, and possible future directions. The final section provides python code which will allow the reader to carry out their own verification of the computations used in our main example.
2. Proofs of Theorem 1.2 and Corollaries 1.3 and 1.5
@@
The two summands which make up our example in Theorem 1.2 are $K=7_{1}$ and its mirror image $\overline{K}$. $K$ is the $(2,7)$-torus knot. A routine calculation establishes that $K$ has signature $\sigma(K)=6$; from the inequality $|\sigma(K)|/2\leq u(K)$ [29] we know that $K$ has unknotting number at least 3. Since any 3 crossing changes in its standard 7-crossing diagram results in the unknot, we have $u(K)=3$. Alternatively, one can appeal to the important and more far-reaching result of Kronheimer and Mrowka [25] that the unknotting number of the $(p,q)$-torus knot $T(p,q)$ is (p1)(q1)/2(p-1)(q-1)/2. It is a straightforward result that unknotting number is preserved under mirror image - hold a mirror up to the unknotting sequences - and so $u(\overline{K})=3$, as well. Consequently, $u(K)+u(\overline{K})=3+3=6$.
@@
We will demonstrate, however, that $L=K\#\overline{K}$ has unknotting number at most $5$. We do this by giving a diagram of $L$ for which changing two crossings gives (a diagram for) the knot $K_{1}=K14a18636$, in the notation from SnapPy. We then show that the knot $K_{2}=K15n81556$ has unknotting number (at most) 2, and has a diagram for which one crossing change yields a diagram for $K_{1}$. Consequently, $L=K\#\overline{K}$ has a diagram for which 2 crossing changes can take us to a diagram of $K_{1}$; an isotopy then takes us to another diagram for $K_{1}$ that admits a crossing change to a diagram of $K_{2}$, which has unknotting number at most $2$. So it takes at most $2+1+2=5$ crossing changes, with isotopies, to transform our diagram of $K\#\overline{K}$ to the unknot. This unknotting sequence therefore has length $5$ (that is, $6$ knots and $5$ crossing changes); consequently, $u(K\#\overline{K})\leq 5$.
@@
The reader can verify in SnapPy/Sage [14],[37] that this is a braid representative for $L$ by using the command K=Link(braid_closure=brd) or K=snappy.Link(braid_closure=brd) (calling the above tuple ‘brd’) and employing the simplify() and deconnect_sum() commands to recover the summands $7_{1}$ and $\overline{7_{1}}$ . Alternatively, one can start from a diagram for $L=7_{1}\#\overline{7_{1}}$, on the left side of Figure 2, and arrive at the diagram of the knot that SnapPy will produce from the braid word above, after a few Reidemeister moves, as indicated in Figure 2. Strands that are not fixed in each step of this isotopy, and in the isotopies that are represented in the diagrams that follow, are represented as thicker (and red) arcs.
@@
Crossing changes on this diagram can be carried out by changing the signs of entries in the tuple, to create a new braid word representing the knot after crossing change. In the above braid word we have underlined the entries relevant to our construction. Following the convention that SnapPy/python starts labeling entries in a tuple at ‘0’, changing the crossings at indices 0 and 1 yields the braid word
$\underline{-1},\underline{4},2,3,3,3,2,3,2,2,4,-3,-3,-3,-3,-1,-3,-2,-3,-3]$
which SnapPy identifies as belonging to the knot $K_{1}=K14a18636$. These crossing changes, and subsequent crossing changes in further diagrams, are marked with a shaded (and yellow) circle in Figure 2. A diagram for this knot $K_{1}$ built from the closure of the braid represented by this braid word is given in Figure 3; the two crossing changes made to obtain it from the rightmost diagram in Figure 2 are circled. We also provide a schematic of an isotopy from this 20-crossing projection to a 14-crossing projection of $K_{1}$. Since unknotting number changes by at most one under crossing change, and we can recover $L$ from $K_{1}$ by reversing these $2$ crossing changes, we have $u(L)\leq u(K_{1})+2$ .
@@
We now demonstrate that $K_{1}$ has unknotting number at most $3$. We do this by making use of a pair of 15-crossing diagrams of the knot $K_{2}=K15n81556$. For this discussion we utilize Dowker-Thistlethwaite (DT) codes [15] for the knots, where, again, each entry of the code represents a crossing in the diagram, and a crossing change can be achieved by changing the sign of an entry of the code. The knot $K_{2}$ has, from the authors’ earlier work on the Bernhard-Jablan conjecture [11], 100 distinct 15-crossing projections. For one of them, with DT code
$\underline{-4},-16,24,26,18,20,28,22,-2,10,12,30,6,8,14]$,
@@
Figure 4. $K14a18636$ to $K15n81556$ (a)
On the other hand, a different diagram for $K_{2}$ is given by the DT code
$4,12,-24,14,18,2,\underline{20},26,8,10,-28,-30,16,-6,-22]$,
A sequence of isotopies interpolating between the two diagrams for $K_{2}$ is shown in Figure 5; a different such sequence can be found in [49]. In the figure, ‘planar isotopy’ means that there are no changes in the combinatorics of the underlying knot diagram; the highlighted arc is straightened to the corresponding arc marked.
Figure 5. $K15n81556$ (a) to $K15n81556$ (b)
When we change a single crossing, at index 6, this yields the DT code
$4,12,-24,14,18,2,\underline{-20},26,8,10,-28,-30,16,-6,-22]$,
which SnapPy identifies as belonging to the knot $K_{3}=K12n412$. The knot $K_{3}$ has unknotting number 1 (as noted, for example, on the Knotinfo site [27]); and in fact changing the crossing in the DT code above for $K_{3}$, at index 13, yields the code
$4,12,-24,14,18,2,-20,26,8,10,-28,-30,16,\underline{6},-22]$,
which is a diagram of the unknot. A sequence of isotopies from the last diagram for $K15n81556$ in Figure 5 to the unknot, after making these two crossing changes, is shown in Figure 6. Consequently, we can obtain the unknot starting from $K_{1}=K14a18636$ with three crossing changes; first to $K_{2}=K15n81556$, then to $K_{3}=K12n412$, and finally to the unknot.
Figure 6. $K15n81556$ (b) to the unknot
@@
It is a well-known fact that in computing unknotting number, crossing changes and isotopies can, in a certain sense, be exchanged, at the expense of the crossing changes taking place in a more complicated diagram. That is, every knot $K$ has a diagram $D$ so that $u(K)$ crossing changes in that diagram yields a diagram for the unknot. After the authors posted an earlier version of this paper to arxiv.org, using some of the sequences of crossing changes and isotopies that we found above Wang and Zhang [49] found such a diagram $D$ for $7_{1}\#\overline{7_{1}}$ for which $5$ crossing changes in the diagram yields a diagram of the unknot. Their diagram for $7_{1}\#\overline{7_{1}}$ has $56$ crossings.
Next we turn our attention to the proof of Corollary 1.3, extending our main example to an infinite class of examples. A knot $K_{1}$ is called Gordian adjacent to the knot $K_{2}$ [16] if $u(K_{2})>u(K_{1})$ and there are $u(K_{2})-u(K_{1})$ crossing changes on a diagram for $K_{2}$ that results in a diagram for $K_{1}$. Put differently, $K_{2}$ possesses a minimal unknotting sequence which contains the knot $K_{1}$. For any pair of knots $L_{1},L_{2}$ that the knot $7_{1}$ is Gordian adjacent to, we can build an unknotting sequence for $L_{1}\#\overline{L_{2}}$ using fewer than $u(L_{1})+u(L_{2})$ crossing changes. In particular, starting with $L_{1}\#\overline{L_{2}}$, applying $u(L_{1})-3=u(L_{1})-u(7_{1})$ crossing changes to the $L_{1}$-summand, and $u(L_{2})-3$ crossing changes to the $\overline{L_{2}}$ summand will yield a diagram for $7_{1}\#\overline{7_{1}}$, and this knot can be unknotted with $5$ crossing changes (after isotopy). Consequently, $(u(L_{1})-3)+(u(L_{2})-3)+5=u(L_{1})+u(L_{2})-1$ crossing changes suffice to unknot $L_{1}\#\overline{L_{2}}$, and so $u(L_{1}\#\overline{L_{2}})\leq u(L_{1})+u(L_{2})-1<u(L_{1})+u(L_{2})$.
One family of knots which can quickly be seen to satisfy this condition is the set of torus knots $T(2,2k+1)$ with $k\geq 3$. These knots have unknotting number (21)((2k+1)1)/2=k(2-1)((2k+1)-1)/2=k, and changing any k3k-3 crossings in their standard diagrams will yield $T(2,7)=7_{1}$ (since changing a crossing in the standard diagram for $T(2,N)$ will, after a Reidemeister II move, yield the standard diagram for T(2,N2)T(2,N-2), and we are then done by induction). This yields Corollary 1.3. $\square$
Further extensions of Theorem 1.2 can be made to pbtain more examples in a similar way. Several authors have explored the construction of minimal unknotting sequences for torus knots [7],[43], and the problem of Gordian adjacency for torus knots [16],[10]. Feller’s work [16] shows that all torus knots $T(p,q)$ except (possibly) $T(2,3),T(2,5),T(3,4),T(4,5)$ and $T(5,6)$ have $7_{1}=T(2,7)$ Gordian adjacent to them. Tetsuya Abe [1] has pointed out that work of Kanenobu [21] shows that $T(2,7)$ is Gordian adjacent to both $T(4,5)$ and $T(5,6)$, and has provided the authors with explicit unknotting sequences to show this. $T(2,7)$ cannot be Gordian adjacent to the knots $T(2,3),T(2,5)$ and $T(3,4)$, however, because their unknotting numbers are too low; they have unknotting numbers 1, 2, and 3, respectively. Taken together, these results show that nearly every torus knot can be used to build a counterexample to the additivity conjecture.
@@
For the collection of torus knots
${\mathcal{T}}=\{T(p,q)$ : $2\leq p<q,\ \gcd(p,q)=1,\ (p,q)\not\in\{(2,3),(2,5),(3,4)\}\}$,
any pair of knots $K_{1},K_{2}\in{\mathcal{T}}$ satisfy $u(K_{1}\#\overline{K_{2}})<u(K_{1})+u(K_{2})$.
It is an interesting open question to decide if the torus knots $T(2,3)$, $T(2,5)$ or $T(3,4)$ can be paired with other knots to form a connected sum which also fails additivity.
For an example that is not a torus knot, we can use the knot $10_{139}$ (which, in fact, was a summand of the actual first example that the authors found in proving Theorem 1.2; see Section 3 below). The knot $L_{1}=10_{139}$ is a non-alternating hyperbolic knot; it was shown by Kawamura [22] that $u(L_{1})=4$. In fact, Kawamura showed, even more, that $L_{1}$ has smooth 4-genus $g_{4}(L_{1})=4$ (and $g_{4}(L)\leq u(L)$ for every knot $L$ [29]); the 4-genus of $L_{1}$ was also established by different means (using their tau invariant) by Ozsváth and Szabó [34]. One DT code for $10_{139}$ is given by
@@
which belongs to the knot $7_{1}$. As a result, an unknotting sequence for $7_{1}$ (with $3$ crossing changes) can be augmented with this crossing change to yield an unknotting sequence for $10_{139}$ with $4$ crossing changes, that is, a minimal unknotting sequence. Consequently, $T(2,7)$ is Gordian adjacent to $10_{139}$, yielding: